A certain amount of cahrge Q is to be divided into two parts q and Q-q. These charges are placed at a certain distance so that there is maximum repulsion. How is Q and q related?

Grade:12

1 Answers

Arun
5 years ago
Dear student
 

let parts are sepereted at a distance of d

so repulsion force F= k q(Q-q)/d2

here Q and d are constant so for maximum force differentiate the force w.r.t q

dF/dq = k/d[Q -2q]

for maximum force dF/dq=0

Q-2q=0

Q/q =2

 

Regards

Arun (askIITians forum expert)